OpenStudy (anonymous):

Graph the function. How is the graph a translation of f(x) = x2? y = (x –1)2 + 3

OpenStudy (anonymous):

let y=f(x) x^2=(x-1)2+3 x^2=2x-2+3 x^2=2x+1 x^2-2x-1=0